OpenStudy (anonymous):

Simplify and solve a^1/2/a^2

OpenStudy (anonymous):

\[a ^{1/2}a^2\]\[a ^{(1/2-2)}\]\[a ^{-3/2}\]\[\frac{1}{a ^{(3/2)}}\]\[\frac{1}{\sqrt[2]{a^3}}\]\[\frac{1}{a \sqrt[2]{a}}\]\[\frac{1}{a \sqrt{a}}\]\[\frac{1 \times \sqrt{a}}{a \sqrt{a} \times \sqrt{a}}\]\[\frac{\sqrt{a}}{a^2}\]